# Data-Driven Logistic Regression Ensembles With Applications in Genomics

### v5

@@ 8.11 Results
Lasso 1.92 1.54 2.45 1.61 2.25 1.64 2.15 1.64
EN 1.15 1.07 1.55 1.07 1.42 1.11 1.31 1.10
Adaptive 3.08 3.82 4.18 4.48 4.08 4.70 4.08 5.21
Relaxed 2.46 3.84 3.00 3.51 3.00 3.71 2.69 3.65
MCP 3.38 3.88 3.91 3.78 3.42 3.73 3.15 3.65
SIS-SCAD 3.23 4.41 3.64 4.20 3.25 4.25 2.92 4.14
Split-Lasso-CV 1.46 1.00 1.73 1.01 1.67 1.04 1.31 1.06
Split-EN-CV 1.23 1.03 1.64 1.00 1.33 1.00 1.31 1.00
RGLM-CV 2.00 1.76 2.27 1.76 1.83 1.94 1.92 2.00
RF-CV 1.00 3.88 1.00 3.77 1.00 3.77 1.00 3.75
XGB 3.85 5.22 4.55 5.09 4.17 5.12 3.85 5.08
Table 46: MR and TL relative performances for GSE23400 (part one) and training proportion 0.35.
@@
Lasso 1.06 1.12 1.06 1.12 1.14 1.17 1.08 1.21
EN 1.02 1.03 1.02 1.04 1.05 1.05 1.02 1.08
Adaptive 1.31 1.53 1.25 1.50 1.08 1.40 1.09 1.46
Relaxed 1.16 3.88 1.18 3.25 1.23 4.13 1.23 3.75
MCP 1.35 1.54 1.40 1.60 1.43 1.68 1.40 1.69
SIS-SCAD 1.50 1.77 1.57 1.88 1.64 1.97 1.62 1.98
Split-Lasso-CV 1.03 1.06 1.00 1.04 1.00 1.00 1.00 1.02
Split-EN-CV 1.00 1.00 1.01 1.00 1.02 1.00 1.01 1.00
RGLM-CV 1.10 1.15 1.09 1.15 1.15 1.20 1.08 1.24
RF-CV 1.05 1.36 1.09 1.45 1.14 1.50 1.10 1.54
XGB 1.74 2.25 1.86 2.43 2.03 3.19 1.98 3.24
Table 47: MR and TL relative performances for GSE23400 (part one) and training proportion 0.5.
@@
Lasso 1.14 1.08 1.12 1.13 1.15 1.14 1.13 1.18
EN 1.07 1.02 1.10 1.05 1.05 1.04 1.04 1.08
Adaptive 1.35 1.43 1.09 1.41 1.04 1.34 1.02 1.44
Relaxed 1.21 3.08 1.23 3.07 1.26 3.02 1.19 3.76
MCP 1.41 1.46 1.38 1.52 1.42 1.61 1.38 1.62
SIS-SCAD 1.42 1.63 1.46 1.77 1.41 1.88 1.41 1.94
Split-Lasso 1.06 1.06 1.00 1.03 1.03 1.02 1.00 1.01
Split-EN 1.00 1.00 1.02 1.00 1.00 1.00 1.03 1.00
RGLM-CV 1.10 1.09 1.02 1.13 1.01 1.18 1.05 1.19
RF-CV 1.12 1.40 1.09 1.47 1.12 1.53 1.11 1.55
XGB 1.48 1.82 1.42 1.91 1.71 2.82 1.76 2.95
Table 48: MR and TL relative performances for GSE23400 (part two) and training proportion 0.35.
@@
Lasso 1.07 1.12 1.06 1.03 1.08 1.07 1.08 1.06
EN 1.07 1.07 1.11 1.00 1.08 1.05 1.07 1.05
Adaptive 1.00 1.16 1.00 1.10 1.00 1.20 1.00 1.33
Relaxed 1.09 2.94 1.13 2.50 1.10 2.61 1.19 2.73
MCP 1.10 1.27 1.10 1.20 1.05 1.26 1.07 1.27
SIS-SCAD 1.11 1.45 1.13 1.44 1.10 1.49 1.13 1.52
Split-Lasso-CV 1.10 1.18 1.08 1.03 1.03 1.00 1.03 1.00
Split-EN-CV 1.06 1.10 1.08 1.00 1.03 1.01 1.01 1.00
RGLM-CV 1.05 1.00 1.13 1.01 1.07 1.09 1.10 1.12
RF-CV 1.10 1.39 1.11 1.34 1.12 1.40 1.10 1.42
XGB 1.64 2.01 1.69 1.92 1.61 2.03 1.77 2.67
Table 49: MR and TL relative performances for GSE23400 (part two) and training proportion 0.5.
@@
Lasso 1.10 1.04 1.03 1.01 1.09 1.03 1.03 1.04
EN 1.20 1.06 1.03 1.00 1.03 1.00 1.04 1.03
Adaptive 1.00 1.19 1.00 1.21 1.05 1.23 1.00 1.35
Relaxed 1.14 2.91 1.14 3.18 1.09 2.89 1.05 2.76
MCP 1.12 1.26 1.04 1.21 1.08 1.24 1.05 1.24
SIS-SCAD 1.08 1.40 1.02 1.40 1.06 1.43 1.02 1.44
Split-Lasso-CV 1.24 1.11 1.06 1.03 1.00 1.02 1.02 1.01
Split-EN-CV 1.24 1.11 1.07 1.03 1.00 1.01 1.02 1.00
RGLM-CV 1.20 1.00 1.16 1.03 1.13 1.06 1.11 1.08
RF-CV 1.23 1.44 1.19 1.41 1.23 1.41 1.18 1.43
XGB 1.45 1.59 1.65 2.39 1.57 2.30 1.67 2.43
Table 50: MR and TL relative performances for GSE5364 (esophageal) and training proportion 0.5.
@@
Lasso 2.47 1.95 2.56 2.21 2.67 2.25 2.78 2.23
EN 1.39 1.17 1.41 1.32 1.42 1.33 1.47 1.31
Adaptive 3.03 2.80 3.68 3.19 3.82 3.07 4.50 3.04
Relaxed 2.67 7.00 2.94 7.57 3.33 9.40 3.19 7.97
MCP 4.83 2.83 5.35 3.24 5.42 3.33 5.62 3.30
SIS-SCAD 4.61 2.82 5.09 3.30 5.27 3.35 5.41 3.29
Split-Lasso-CV 1.19 1.08 1.18 1.22 1.36 1.26 1.44 1.26
Split-EN-CV 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
RGLM-CV 2.33 1.70 2.00 1.84 1.85 1.83 1.97 1.81
RF-CV 1.08 2.15 1.18 2.46 1.12 2.48 1.25 2.43
XGB 4.03 2.95 4.26 3.37 4.33 3.41 4.53 3.35
Table 51: MR and TL relative performances for GSE25869 and training proportion 0.5.
@@
Lasso 1.13 1.42 1.24 1.38 1.24 1.34 1.24 1.37
EN 1.10 1.27 1.15 1.29 1.15 1.26 1.17 1.29
Adaptive 1.22 1.24 1.32 1.31 1.33 1.39 1.33 1.36
Relaxed 1.20 2.73 1.24 3.36 1.29 2.22 1.28 2.44
MCP 1.37 1.34 1.49 1.40 1.49 1.42 1.45 1.51
SIS-SCAD 1.40 1.31 1.51 1.40 1.50 1.45 1.51 1.49
Split-Lasso-CV 1.05 1.23 1.08 1.23 1.03 1.19 1.06 1.18
Split-EN-CV 1.05 1.11 1.03 1.12 1.07 1.12 1.02 1.11
RGLM-CV 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
RF-CV 1.08 1.13 1.13 1.20 1.13 1.23 1.11 1.23
XGB 1.33 1.38 1.51 1.57 1.45 1.56 1.52 1.61
Table 52: MR and TL relative performances for GSE10245 and training proportion 0.35.
@@
Lasso 1.21 1.31 1.17 1.33 1.24 1.37 1.27 1.40
EN 1.07 1.11 1.03 1.10 1.05 1.11 1.01 1.15
Adaptive 1.76 1.68 1.59 1.69 1.86 1.81 1.86 1.76
Relaxed 1.36 3.57 1.25 3.14 1.34 3.67 1.32 2.80
MCP 1.99 1.71 1.92 2.44 1.99 1.66 2.02 1.75
SIS-SCAD 2.07 1.70 2.01 1.72 2.06 1.69 2.04 1.69
Split-Lasso-CV 1.06 1.10 1.01 1.05 1.06 1.05 1.00 1.09
Split-EN-CV 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
RGLM-CV 1.39 1.59 1.09 1.30 1.06 1.31 1.08 1.34
RF-CV 1.46 1.50 1.50 1.54 1.54 1.55 1.56 1.57
XGB 2.50 1.95 2.51 2.03 2.63 2.03 2.60 2.38
Table 53: MR and TL relative performances for GSE10245 and training proportion 0.5.
@@
Lasso 1.29 1.31 1.29 1.37 1.36 1.33 1.29 1.35
EN 1.09 1.07 1.11 1.12 1.17 1.08 1.07 1.08
Adaptive 1.72 1.61 2.04 1.83 1.81 1.72 1.85 1.84
Relaxed 1.25 3.00 1.40 4.07 1.49 3.30 1.37 3.67
MCP 2.08 1.67 2.30 1.95 2.50 2.01 2.28 2.03
SIS-SCAD 2.20 1.75 2.45 2.09 2.54 2.13 2.38 2.18
Split-Lasso-CV 1.00 1.06 1.00 1.04 1.00 1.02 1.02 1.05
Split-EN-CV 1.02 1.00 1.01 1.00 1.02 1.00 1.00 1.00
RGLM-CV 1.36 1.84 1.24 1.52 1.26 1.54 1.13 1.55
RF-CV 1.33 1.40 1.51 1.58 1.63 1.61 1.49 1.66
XGB 2.61 2.17 2.80 2.51 3.10 2.77 2.92 2.86
Table 54: MR and TL relative performances for GSE5364 (lung) and training proportion 0.5.
@@
Lasso 2.34 1.68 2.04 1.60 2.36 1.57 2.31 1.58
EN 1.75 1.23 1.45 1.20 1.51 1.12 1.57 1.13
Adaptive 3.25 1.87 3.07 1.92 2.92 1.78 3.06 1.74
Relaxed 2.57 5.31 2.29 4.60 2.51 4.32 2.59 4.64
MCP 3.47 1.98 3.36 2.05 3.53 1.98 3.74 2.02
SIS-SCAD 3.35 1.87 3.36 1.95 3.60 1.89 3.65 1.88
Split-Lasso-CV 1.00 1.05 1.16 1.07 1.17 1.03 1.13 1.03
Split-EN-CV 1.04 1.00 1.00 1.00 1.08 1.00 1.13 1.00
RGLM-CV 1.63 1.47 1.16 1.38 1.02 1.25 1.09 1.24
RF-CV 1.12 1.50 1.00 1.51 1.00 1.42 1.00 1.41
XGB 3.14 1.87 2.98 1.87 3.09 1.77 3.37 2.09
Table 55: MR and TL relative performances for GSE5364 (thyroid) and training proportion 0.5.
@@
Lasso 1.39 1.24 1.43 1.22 1.47 1.36 1.35 1.29
EN 1.16 1.10 1.28 1.10 1.25 1.21 1.18 1.21
Adaptive 1.22 1.08 1.43 1.14 1.79 1.33 2.00 1.37
Relaxed 1.57 3.07 1.66 3.38 1.57 2.85 1.56 2.88
MCP 1.70 1.22 1.83 1.25 1.98 1.36 1.88 1.34
SIS-SCAD 1.84 1.20 1.97 1.21 1.98 1.32 1.92 1.30
Split-Lasso-CV 1.03 1.13 1.05 1.03 1.04 1.12 1.08 1.14
Split-EN-CV 1.00 1.00 1.00 1.00 1.00 1.08 1.02 1.06
RGLM-CV 1.11 1.08 1.08 1.04 1.03 1.00 1.00 1.00
RF-CV 1.01 1.03 1.03 1.04 1.12 1.15 1.02 1.13
XGB 1.94 1.21 2.18 1.23 2.21 1.61 2.07 1.55
Table 56: MR and TL relative performances for GSE21942 and training proportion 0.5.
@@
Lasso 1.82 1.32 1.50 1.25 1.75 1.36 1.75 1.40
EN 1.00 1.00 1.00 1.00 1.00 1.01 1.00 1.05
Adaptive 4.91 2.69 4.75 2.86 5.08 3.07 4.42 3.02
Relaxed 2.64 3.20 2.17 2.61 2.67 3.14 2.42 2.39
MCP 7.73 4.23 7.17 4.09 7.08 3.85 6.50 3.55
SIS-SCAD 7.73 4.31 6.75 4.14 6.83 3.90 6.75 3.78
Split-Lasso-CV 2.18 1.18 1.33 1.10 1.17 1.00 1.25 1.02
Split-EN-CV 2.00 1.18 1.92 1.12 1.75 1.04 1.58 1.00
RGLM-CV 2.45 2.58 2.33 2.55 2.92 2.47 3.25 2.40
RF-CV 4.09 3.75 3.75 3.67 3.67 3.45 3.58 3.34
XGB 9.45 5.23 8.67 5.09 8.58 4.76 5.67 5.98
Table 57: MR and TL relative performances for GSE14905 and training proportion 0.5.
@@
Lasso 1.82 1.65 1.78 1.67 1.72 1.60 2.00 1.61
EN 1.32 1.11 1.30 1.12 1.53 1.12 1.64 1.13
Adaptive 2.13 2.48 2.11 2.34 1.83 2.26 1.91 2.36
Relaxed 1.79 5.06 1.97 4.50 1.72 4.44 1.91 3.79
MCP 2.74 3.26 2.81 3.37 2.86 3.25 3.12 3.19
SIS-SCAD 2.58 3.58 2.41 3.50 2.44 3.39 2.58 3.26
Split-Lasso-CV 1.34 1.04 1.30 1.10 1.47 1.06 1.85 1.10
Split-EN-CV 1.18 1.00 1.14 1.00 1.22 1.00 1.52 1.00
RGLM-CV 1.95 1.62 1.65 1.47 1.67 1.39 1.85 1.41
RF-CV 1.00 2.59 1.00 2.63 1.00 2.56 1.00 2.53
XGB 2.39 3.47 2.46 3.53 2.44 7.49 2.67 7.33

### v6

@@ 8.11 Results
Split-Lasso-CV 2.31 1.00 2.31 1.05 2.31 1.07 2.31 1.08
Split-EN-CV 2.31 1.03 2.31 1.00 2.31 1.00 2.31 1.00
Lasso 2.26 1.92 2.26 1.95 2.26 1.89 2.26 1.89
EN 2.31 1.12 2.31 1.13 2.31 1.15 2.31 1.15
Adaptive 2.24 3.92 2.22 4.64 2.21 4.98 2.21 5.10
Relaxed 2.25 5.36 2.24 5.27 2.26 5.38 2.25 4.61
MCP 2.21 3.99 2.21 3.94 2.20 3.97 2.21 3.86
SIS-SCAD 2.20 4.66 2.21 4.64 2.21 4.52 2.21 4.43
RuleFit 2.20 7.17 2.20 7.09 2.20 7.08 2.20 7.01
RE-Lasso-100 2.30 1.17 2.29 1.23 2.30 1.25 2.30 1.34
RE-EN-100 2.31 1.01 2.31 1.04 2.31 1.08 2.31 1.11
RGLM-100 2.29 2.00 2.30 1.92 2.29 2.07 2.29 2.06
RF-500 1.00 14.32 1.00 14.14 1.00 14.15 1.00 13.99
XGB 2.20 5.27 2.20 5.24 2.20 5.22 2.20 5.17
Table 74: ACC and TSL relative performances for GSE23400 (part one) and training proportion 0.35.
@@
Split-Lasso-CV 1.05 1.12 1.06 1.08 1.06 1.04 1.06 1.05
Split-EN-CV 1.05 1.05 1.05 1.07 1.06 1.05 1.06 1.03
Lasso 1.04 1.20 1.05 1.20 1.05 1.21 1.05 1.18
EN 1.05 1.07 1.05 1.09 1.06 1.09 1.06 1.09
Adaptive 1.01 1.71 1.03 1.73 1.04 1.61 1.04 1.72
Relaxed 1.03 3.82 1.04 3.65 1.04 3.66 1.04 4.15
MCP 1.02 1.59 1.02 1.66 1.02 1.68 1.02 1.67
SIS-SCAD 1.01 1.80 1.01 1.95 1.01 2.00 1.01 2.00
RuleFit 1.00 3.28 1.00 3.59 1.00 3.62 1.00 3.64
RE-Lasso-100 1.05 1.01 1.06 1.01 1.06 1.02 1.06 1.01
RE-EN-100 1.05 1.00 1.06 1.00 1.06 1.00 1.06 1.00
RGLM-100 1.04 1.23 1.05 1.21 1.05 1.22 1.06 1.25
RF-500 1.04 1.30 1.05 1.41 1.05 1.46 1.05 1.49
XGB 1.00 2.18 1.00 2.37 1.00 2.43 1.00 2.45
Table 75: ACC and TSL relative performances for GSE23400 (part one) and training proportion 0.5.
@@
Split-Lasso-CV 1.03 1.11 1.03 1.08 1.03 1.03 1.03 1.03
Split-EN-CV 1.03 1.06 1.03 1.07 1.03 1.04 1.02 1.02
Lasso 1.02 1.13 1.02 1.15 1.02 1.16 1.02 1.16
EN 1.03 1.06 1.02 1.07 1.02 1.06 1.02 1.07
Adaptive 1.00 1.56 1.01 1.57 1.01 1.62 1.01 1.66
Relaxed 1.02 2.57 1.02 2.23 1.02 2.20 1.02 2.76
MCP 1.00 1.55 1.01 1.59 1.01 1.60 1.00 1.61
SIS-SCAD 1.01 1.72 1.01 1.78 1.01 1.80 1.01 1.82
RuleFit 1.00 2.13 1.00 2.36 1.00 2.37 1.00 2.34
RE-Lasso-100 1.03 1.00 1.03 1.00 1.03 1.02 1.03 1.02
RE-EN-100 1.03 1.03 1.03 1.01 1.03 1.00 1.03 1.00
RGLM-100 1.03 1.13 1.03 1.19 1.03 1.18 1.03 1.18
RF-500 1.03 1.33 1.02 1.44 1.02 1.44 1.02 1.45
XGB 1.01 1.78 1.00 1.91 1.00 1.96 1.00 1.98
Table 76: ACC and TSL relative performances for GSE23400 (part two) and training proportion 0.35.
@@
Split-Lasso-CV 1.04 1.21 1.05 1.11 1.04 1.06 1.04 1.05
Split-EN-CV 1.04 1.14 1.04 1.10 1.05 1.08 1.04 1.05
Lasso 1.04 1.13 1.05 1.11 1.05 1.13 1.04 1.11
EN 1.05 1.09 1.04 1.08 1.04 1.07 1.04 1.07
Adaptive 1.04 1.30 1.05 1.21 1.05 1.27 1.04 1.25
Relaxed 1.03 3.27 1.04 2.79 1.04 3.14 1.04 3.13
MCP 1.04 1.32 1.04 1.31 1.04 1.33 1.03 1.32
SIS-SCAD 1.03 1.58 1.03 1.60 1.03 1.62 1.03 1.60
RuleFit 1.00 2.92 1.01 2.96 1.00 3.02 1.00 2.99
RE-Lasso-100 1.05 1.00 1.05 1.00 1.05 1.00 1.05 1.00
RE-EN-100 1.05 1.04 1.05 1.01 1.05 1.00 1.05 1.04
RGLM-100 1.04 1.03 1.05 1.09 1.04 1.14 1.04 1.14
RF-500 1.03 1.30 1.04 1.34 1.04 1.41 1.04 1.46
XGB 1.00 2.00 1.00 2.05 1.00 2.04 1.00 1.97
Table 77: ACC and TSL relative performances for GSE23400 (part two) and training proportion 0.5.
@@
Split-Lasso-CV 1.02 1.10 1.03 1.10 1.04 1.07 1.03 1.02
Split-EN-CV 1.03 1.07 1.03 1.07 1.03 1.06 1.03 1.03
Lasso 1.03 1.12 1.02 1.14 1.03 1.15 1.03 1.14
EN 1.03 1.08 1.03 1.04 1.03 1.08 1.03 1.05
Adaptive 1.04 1.23 1.03 1.23 1.04 1.28 1.03 1.24
Relaxed 1.03 2.83 1.03 2.98 1.03 3.41 1.02 2.82
MCP 1.04 1.21 1.03 1.27 1.04 1.28 1.03 1.28
SIS-SCAD 1.04 1.34 1.04 1.42 1.04 1.48 1.04 1.50
RuleFit 1.00 1.94 1.00 2.09 1.00 2.02 1.00 2.02
RE-Lasso-100 1.04 1.01 1.03 1.00 1.04 1.00 1.04 1.04
RE-EN-100 1.03 1.03 1.04 1.01 1.04 1.02 1.04 1.00
RGLM-100 1.02 1.00 1.03 1.06 1.03 1.08 1.03 1.10
RF-500 1.02 1.26 1.02 1.31 1.03 1.38 1.02 1.43
XGB 1.00 1.59 1.00 1.58 1.00 1.70 1.00 1.60
Table 78: ACC and TSL relative performances for GSE5364 (Esophageal) and training proportion 0.5.
@@
Split-Lasso-CV 2.12 1.06 2.13 1.20 2.19 1.24 2.15 1.32
Split-EN-CV 2.14 1.01 2.15 1.00 2.22 1.02 2.18 1.08
Lasso 2.00 1.70 2.01 1.84 2.06 1.96 2.03 2.10
EN 2.10 1.20 2.10 1.29 2.16 1.32 2.13 1.35
Adaptive 1.92 2.59 1.86 3.00 1.91 3.00 1.75 3.65
Relaxed 1.98 5.60 1.97 6.49 2.03 6.53 2.01 6.69
MCP 1.79 2.72 1.78 3.13 1.80 3.34 1.81 3.35
SIS-SCAD 1.81 2.68 1.80 3.05 1.85 3.20 1.83 3.25
RuleFit 1.82 6.36 1.82 7.05 1.87 7.39 1.85 7.54
RE-Lasso-100 2.11 1.16 2.12 1.17 2.17 1.18 2.16 1.23
RE-EN-100 2.14 1.00 2.15 1.02 2.21 1.00 2.20 1.00
RGLM-100 2.02 1.61 2.05 1.69 2.15 1.73 2.13 1.80
RF-500 1.00 4.55 1.00 5.04 1.00 5.27 1.00 5.38
XGB 1.84 2.85 1.84 3.17 1.89 3.32 1.87 3.38
Table 79: ACC and TSL relative performances for GSE25869 and training proportion 0.35.
@@
Split-Lasso-CV 1.10 1.66 1.18 1.43 1.23 1.29 1.24 1.21
Split-EN-CV 1.10 1.51 1.19 1.34 1.23 1.21 1.25 1.15
Lasso 1.07 1.59 1.14 1.47 1.20 1.36 1.18 1.32
EN 1.09 1.50 1.17 1.36 1.22 1.28 1.22 1.23
Adaptive 1.00 1.29 1.00 1.36 1.00 1.41 1.00 1.45
Relaxed 1.07 4.05 1.14 3.40 1.19 3.37 1.17 3.48
MCP 1.04 1.39 1.06 1.46 1.09 1.45 1.09 1.46
SIS-SCAD 1.03 1.35 1.10 1.40 1.14 1.39 1.14 1.41
RuleFit 1.06 2.26 1.13 2.36 1.17 2.40 1.17 2.43
RE-Lasso-100 1.09 1.30 1.17 1.22 1.22 1.17 1.22 1.16
RE-EN-100 1.10 1.38 1.18 1.23 1.23 1.17 1.24 1.14
RGLM-100 1.08 1.00 1.17 1.00 1.20 1.00 1.21 1.00
RF-500 1.08 1.05 1.14 1.11 1.19 1.12 1.20 1.14
XGB 1.06 1.32 1.12 1.39 1.16 1.39 1.16 1.40
Table 80: ACC and TSL relative performances for GSE25869 and training proportion 0.5.
@@
Split-Lasso-CV 1.09 1.46 1.07 1.32 1.14 1.30 1.19 1.23
Split-EN-CV 1.08 1.34 1.07 1.30 1.14 1.23 1.20 1.23
Lasso 1.06 1.48 1.04 1.51 1.10 1.51 1.16 1.47
EN 1.08 1.38 1.06 1.33 1.12 1.33 1.17 1.33
Adaptive 1.00 1.27 1.00 1.35 1.00 1.51 1.00 1.59
Relaxed 1.08 3.39 1.03 3.59 1.09 3.70 1.14 3.31
MCP 1.04 1.35 1.02 1.42 1.08 1.51 1.10 1.49
SIS-SCAD 1.04 1.30 1.00 1.38 1.06 1.44 1.10 1.45
RuleFit 1.04 1.34 1.01 1.43 1.07 1.48 1.12 1.51
RE-Lasso-100 1.08 1.24 1.05 1.19 1.11 1.24 1.18 1.23
RE-EN-100 1.08 1.24 1.06 1.20 1.14 1.20 1.17 1.20
RGLM-100 1.09 1.02 1.06 1.00 1.13 1.00 1.17 1.00
RF-500 1.08 1.00 1.06 1.08 1.12 1.13 1.17 1.15
XGB 1.02 1.21 1.01 1.29 1.06 1.35 1.12 1.36
Table 81: ACC and TSL relative performances for GSE10245 and training proportion 0.35.
@@
Split-Lasso-CV 1.14 1.12 1.15 1.05 1.15 1.05 1.14 1.07
Split-EN-CV 1.14 1.10 1.15 1.00 1.15 1.00 1.15 1.00
Lasso 1.12 1.43 1.12 1.47 1.12 1.50 1.11 1.46
EN 1.14 1.16 1.14 1.17 1.14 1.20 1.14 1.16
Adaptive 1.06 1.76 1.06 1.86 1.04 1.97 1.03 1.93
Relaxed 1.10 4.00 1.11 3.66 1.11 3.34 1.11 3.66
MCP 1.05 1.78 1.05 1.75 1.05 1.84 1.04 1.74
SIS-SCAD 1.05 1.78 1.05 1.74 1.05 1.79 1.05 1.71
RuleFit 1.00 4.38 1.00 4.39 1.00 4.42 1.01 4.22
RE-Lasso-100 1.14 1.10 1.15 1.09 1.14 1.23 1.13 1.18
RE-EN-100 1.15 1.00 1.15 1.00 1.15 1.01 1.14 1.05
RGLM-100 1.11 1.72 1.12 1.49 1.13 1.37 1.13 1.32
RF-500 1.11 1.48 1.10 1.51 1.10 1.54 1.10 1.50
XGB 1.00 1.99 1.00 1.99 1.00 1.99 1.00 1.90
Table 82: ACC and TSL relative performances for GSE10245 and training proportion 0.5.
@@
Split-Lasso-CV 1.14 1.13 1.15 1.04 1.15 1.06 1.15 1.05
Split-EN-CV 1.14 1.03 1.15 1.00 1.15 1.00 1.15 1.00
Lasso 1.12 1.36 1.11 1.43 1.12 1.44 1.11 1.50
EN 1.14 1.11 1.14 1.15 1.14 1.13 1.15 1.14
Adaptive 1.05 1.93 1.03 2.31 1.01 2.31 1.00 2.32
Relaxed 1.11 3.27 1.11 2.98 1.12 2.61 1.11 3.06
MCP 1.04 1.89 1.05 2.02 1.06 1.98 1.07 1.91
SIS-SCAD 1.02 2.02 1.03 2.22 1.05 2.16 1.05 2.06
RuleFit 1.00 4.32 1.00 4.75 1.01 4.73 1.01 4.55
RE-Lasso-100 1.14 1.09 1.14 1.15 1.13 1.29 1.12 1.26
RE-EN-100 1.14 1.00 1.14 1.07 1.13 1.11 1.14 1.12
RGLM-100 1.09 2.08 1.12 1.61 1.13 1.46 1.13 1.42
RF-500 1.10 1.57 1.11 1.67 1.10 1.68 1.12 1.63
XGB 1.00 2.32 1.00 2.54 1.00 2.56 1.01 2.47
Table 83: ACC and TSL relative performances for GSE5364 (Lung) and training proportion 0.5.
@@
Split-Lasso-CV 1.76 1.03 1.75 1.05 1.76 1.01 1.74 1.01
Split-EN-CV 1.76 1.00 1.76 1.00 1.75 1.00 1.77 1.00
Lasso 1.57 1.64 1.57 1.62 1.58 1.59 1.60 1.56
EN 1.70 1.21 1.68 1.19 1.69 1.16 1.70 1.12
Adaptive 1.52 1.75 1.48 1.83 1.49 1.81 1.44 1.80
Relaxed 1.54 4.99 1.54 4.33 1.54 4.25 1.54 5.00
MCP 1.47 1.92 1.45 1.95 1.43 1.94 1.41 1.96
SIS-SCAD 1.44 1.86 1.44 1.87 1.44 1.83 1.43 1.83
RuleFit 1.44 4.32 1.47 4.34 1.45 4.27 1.46 4.14
RE-Lasso-100 1.73 1.05 1.73 1.07 1.74 1.07 1.74 1.07
RE-EN-100 1.74 1.01 1.74 1.01 1.74 1.00 1.75 1.00
RGLM-100 1.69 1.34 1.74 1.33 1.75 1.29 1.75 1.31
RF-500 1.00 2.69 1.00 2.69 1.00 2.62 1.00 2.58
XGB 1.48 1.87 1.48 1.85 1.48 1.83 1.48 1.79
Table 84: ACC and TSL relative performances for GSE5364 (Thyroid) and training proportion 0.5.
@@
Split-Lasso-CV 1.09 1.23 1.12 1.12 1.11 1.10 1.10 1.08
Split-EN-CV 1.09 1.11 1.12 1.05 1.11 1.06 1.10 1.05
Lasso 1.04 1.38 1.05 1.38 1.06 1.43 1.05 1.34
EN 1.07 1.25 1.08 1.22 1.09 1.21 1.08 1.20
Adaptive 1.07 1.20 1.08 1.21 1.05 1.26 1.04 1.25
Relaxed 1.03 3.16 1.04 3.01 1.04 3.08 1.04 2.86
MCP 1.00 1.32 1.02 1.32 1.02 1.30 1.00 1.31
SIS-SCAD 1.00 1.33 1.00 1.34 1.00 1.29 1.00 1.29
RuleFit 1.03 1.99 1.04 1.94 1.05 1.91 1.04 1.90
RE-Lasso-100 1.09 1.09 1.11 1.05 1.11 1.07 1.11 1.08
RE-EN-100 1.10 1.07 1.12 1.03 1.12 1.00 1.11 1.00
RGLM-100 1.08 1.19 1.11 1.10 1.11 1.00 1.12 1.01
RF-500 1.10 1.00 1.12 1.00 1.13 1.01 1.12 1.05
XGB 1.02 1.19 1.03 1.19 1.04 1.15 1.03 1.16
Table 85: ACC and TSL relative performances for GSE21942 and training proportion 0.5.
@@
Split-Lasso-CV 2.23 1.31 2.22 1.09 2.21 1.13 2.20 1.05
Split-EN-CV 2.22 1.30 2.20 1.30 2.20 1.28 2.20 1.19
Lasso 2.23 1.26 2.21 1.25 2.20 1.30 2.19 1.29
EN 2.25 1.00 2.24 1.00 2.23 1.00 2.23 1.00
Adaptive 2.15 2.32 2.10 2.60 2.12 2.52 2.09 2.52
Relaxed 2.19 4.46 2.18 4.40 2.17 4.53 2.17 3.55
MCP 2.06 3.37 2.04 3.32 2.02 3.46 2.02 3.23
SIS-SCAD 2.03 3.61 2.02 3.52 2.02 3.50 2.01 3.29
RuleFit 2.01 8.05 2.00 7.84 1.99 7.80 1.98 7.34
RE-Lasso-100 2.24 1.08 2.21 1.10 2.21 1.16 2.20 1.15
RE-EN-100 2.22 1.22 2.19 1.25 2.19 1.33 2.19 1.35
RGLM-100 2.19 2.07 2.18 2.14 2.19 2.29 2.17 2.20
RF-500 1.00 8.23 1.00 8.03 1.00 7.99 1.00 7.51
XGB 2.01 4.27 2.00 4.18 1.99 4.17 1.98 3.91
Table 86: ACC and TSL relative performances for GSE14905 and training proportion 0.5.
Method MR TL MR TL MR TL MR TL
Split-Lasso-CV 1.04 1.04 1.04 1.02 1.04 1.10 1.04 1.08
Split-EN-CV 1.05 1.01 1.05 1.01 1.04 1.02 1.04 1.05
Lasso 1.02 1.69 1.03 1.73 1.03 1.65 1.03 1.73
EN 1.05 1.08 1.04 1.15 1.04 1.17 1.04 1.20
Adaptive 1.02 2.02 1.02 2.29 1.02 2.34 1.01 2.43
Relaxed 1.02 4.87 1.03 3.69 1.03 4.19 1.02 4.21
MCP 1.02 2.84 1.02 2.86 1.02 2.79 1.02 2.78
SIS-SCAD 1.02 3.24 1.02 3.22 1.02 3.19 1.02 3.14
RuleFit 1.00 5.83 1.00 5.85 1.00 5.78 1.00 5.76
RE-Lasso-100 1.04 1.24 1.03 1.25 1.03 1.25 1.03 1.28
RE-EN-100 1.05 1.00 1.05 1.00 1.04 1.00 1.05 1.00
RGLM-100 1.03 1.66 1.03 1.42 1.03 1.40 1.03 1.41
RF-500 1.05 1.68 1.05 1.86 1.05 1.99 1.06 2.15
XGB 1.00 3.59 1.00 3.61 1.00 3.58 1.00 3.57